# Neural Network Diffusion

### v1

@@ Abstract
Diffusion models have achieved remarkable success in image and video generation. In this work, we demonstrate that diffusion models can also generate high-performing neural network parameters. Our approach is simple, utilizing an autoencoder and a standard latent diffusion model. The autoencoder extracts latent representations of a subset of the trained network parameters. A diffusion model is then trained to synthesize these latent parameter representations from random noise. It then generates new representations that are passed through the autoencoder’s decoder, whose outputs are ready to use as new subsets of network parameters. Across various architectures and datasets, our diffusion process consistently generates models of comparable or improved performance over trained networks, with minimal additional cost. Notably, we empirically find that the generated models perform differently with the trained networks. Our results encourage more exploration on the versatile use of diffusion models.
Machine Learning, ICML
@@ 1 Introduction
The origin of diffusion models can be traced back to non-equilibrium thermodynamics (Jarzynski, 1997; Sohl-Dickstein et al., 2015). Diffusion processes were first utilized to progressively remove noise from inputs and generate clear images in (Sohl-Dickstein et al., 2015). Later works, such as DDPM (Ho et al., 2020) and DDIM (Song et al., 2021), refine diffusion models, with a training paradigm characterized by forward and reverse processes.
@@
Despite the great success of diffusion models in visual generation, their potential in other domains remains relatively underexplored. In this work, we demonstrate the surprising capability of diffusion models in generating high-performing model parameters, a task fundamentally distinct from traditional visual generation. Parameter generation focuses on creating neural network parameters that can perform well on given tasks. It has been explored from prior and probability modeling aspects, i.e. stochastic neural network (Sompolinsky et al., 1988; Bottou et al., 1991; Wong, 1991; Schmidt et al., 1992; Murata et al., 1994) and Bayesian neural network (Neal, 2012; Kingma & Welling, 2013; Rezende et al., 2014; Kingma et al., 2015; Gal & Ghahramani, 2016). However, using a diffusion model in parameter generation has not been well-explored yet.
Figure 1: The top: illustrates the standard diffusion process in image generation. The bottom: denotes the parameter distribution of batch normalization (BN) during the training CIFAR-100 with ResNet-18. The upper half of the bracket: BN weights. The lower half of the bracket: BN biases.
Taking a closer look at the neural network training and diffusion models, the diffusion-based image generation shares commonalities with the stochastic gradient descent (SGD) learning process in the following aspects (illustrated in Fig. 1). i) Both neural network training and the reverse process of diffusion models can be regarded as transitions from random noise/initialization to specific distributions. ii) High-quality images and high-performing parameters can also be degraded into simple distributions, such as Gaussian distribution, through multiple noise additions.
Based on the observations above, we introduce a novel approach for parameter generation, named neural network diffusion (p-diff, p stands for parameter), which employs a standard latent diffusion model to synthesize a new set of parameters. That is motivated by the fact that the diffusion model has the capability to transform a given random distribution to a specific one. Our method is simple, comprising an autoencoder and a standard latent diffusion model to learn the distribution of high-performing parameters. First, for a subset of parameters of models trained by the SGD optimizer, the autoencoder is trained to extract the latent representations for these parameters. Then, we leverage a standard latent diffusion model to synthesize latent representations from random noise. Finally, the synthesized latent representations are passed through the trained autoencoder’s decoder to yield new high-performing model parameters.
@@
Figure 2: Our approach consists of two processes, named parameter autoencoder and generation. Parameter autoencoder aims to extract the latent representations and reconstruct model parameters via the decoder. The extracted representations are used to train a standard latent diffusion model (LDM). In the inference, the random noise is fed into LDM and trained decoder to obtain the generated parameters.
2 Nerual Network Diffusion
@@ 2.1 Preliminaries of diffusion models
Diffusion models typically consist of forward and reverse processes in a multi-step chain indexed by timesteps. We introduce these two processes in the following.
@@
Different from the forward process, the reverse process aims to train a denoising network to recursively remove the noise from $x_{t}$. It moves backward on the multi-step chain as t decreases from $T$ to 0. Mathematically, the reverse process can be formulated as follows,
@@
$L_{\mathrm{dm}}=\mathcal{D}_{KL}(q(x_{t-1}|x_{t},x_{0})||p_{\theta}(x_{t-1}|x_{t})),$ (3)
where the $\mathcal{D}_{KL}(\cdot||\cdot)$ denotes the Kullback–Leibler (KL) divergence that is normally used to compute the difference between two distributions.
@@
The goal of the training diffusion model is to find the reverse transitions that maximize the likelihood of the forward transitions in each time step $t$. In practice, training equivalently consists of minimizing the variational upper bound. The inference procedure aims to generate novel samples from random noise via the optimized denoising parameters $\theta^{*}$ and the multi-step chains in the reverse process.
@@ 2.2 Overview
We propose neural network diffusion (p-diff), which aims to generate high-performing parameters from random noise. As illustrated in Fig. 2, our method consists of two processes, named parameter autoencoder and generation. Given a set of trained high-performing models, we first select a subset of these parameters and flatten them into 1-dimensional vectors. Subsequently, we introduce an encoder to extract latent representations from these vectors, accompanied by a decoder responsible for reconstructing the parameters from latent representations. Then, a standard latent diffusion model is trained to synthesize latent representations from random noise. After training, we utilize p-diff to generate new parameters via the following chain: random noise $\rightarrow$ reverse process $\rightarrow$ trained decoder $\rightarrow$ generated parameters.
@@ 2.3 Parameter autoencoder
In our paper, we default to synthesizing a subset of model parameters. Therefore, to collect the training data for the autoencoder, we train a model from scratch and densely save checkpoints in the last epoch. It is worth noting that we only update the selected subset of parameters via SGD optimizer and fix the remained parameters of the model. The saved subsets of parameters $S=[s_{1},\ldots,s_{k},\ldots,s_{K}]$ is utilized to train the autoencoder, where $K$ is the number of the training samples. For some large architectures that have been trained on large-scale datasets, considering the cost of training them from scratch, we fine-tune a subset of the parameters of the pre-trained model and densely save the fine-tuned parameters as training samples.
@@
$L_{\mathrm{MSE}}=\frac{1}{K}\sum\begin{subarray}{c}K\\ \\ 1\end{subarray}\|v_{k}-v_{k}^{\prime}\|^{2},$ (5)
where $v_{k}^{\prime}$ is the reconstructed parameters of $k$-th model.
Table 1: We present results in the format of ‘original / ensemble / p-diff’. Our method obtains similar or even higher performance than baselines. The results of p-diff is average in three runs. Bold entries are best results.
@@
2.4 Parameter generation
One of the most direct strategies is to synthesize the novel parameters via a diffusion model. However, the memory cost of this operation is too heavy, especially when the dimension of $V$ is ultra-large. Based on this consideration, we apply the diffusion process to the latent representations by default. For $Z=[z_{1}^{0},\cdots,z_{k}^{0},\cdots,z_{K}^{0}]$ extracted from parameter autoencoder, we use the optimization of DDPM (Ho et al., 2020) as follows,
@@ 2.4 Parameter generation
where $t$ is uniform between $1$ and $T$, the sequence of hyperparameters $\overline{\alpha}_{t}$ indicates the noise strength at each step, $\epsilon$ is the added Gaussian noise, $\epsilon_{\theta}(\cdot)$ denotes the denoising network that parameterized by $\theta$. After finishing the training of the parameter generation, we directly fed random noise into the reverse process and the trained decoder to generate a new set of high-performing parameters. These generated parameters are concatenated with the remained model parameters to form new models for evaluation. Neural network parameters and image pixels exhibit significant disparities in several key aspects, including data type, dimensions, range, and physical interpretation. Different from images, neural network parameters mostly have no spatial relevance, so we replace 2D convolutions with 1D convolutions in our parameter autoencoder and parameter generation processes.
3 Experiments
In this section, We first introduce the setup for reproducing. Then, we report the result comparisons and ablation studies.
3.1 Setup
Datasets and architectures.
We evaluate our approach across a wide range of datasets, including MNIST (LeCun et al., 1998), CIFAR-10/100 (Krizhevsky et al., 2009), ImageNet-1K (Deng et al., 2009), STL-10 (Coates et al., 2011), Flowers (Nilsback & Zisserman, 2008), Pets (Parkhi et al., 2012), and F-101 (Bossard et al., 2014) to study the effectiveness of our method. We mainly conduct experiments on ResNet-18/50 (He et al., 2016), ViT-Tiny/Base (Dosovitskiy et al., 2020), and ConvNeXt-T/B (Liu et al., 2022).
Training details.
The autoencoder and latent diffusion model both include a 4-layer 1D CNNs-based encoder and decoder. We default to collecting 200 training data for all architectures. For ResNet-18/50, we train the models from scratch. In the last epoch, we continue to train the last two normalization layers and fix the other parameters. We save 200 checkpoints in the last epoch, i.e., original models. For ViT-Tiny/Base and ConvNeXt-T/B, we fine-tune the last two normalization parameters of the released model in the timm library (Wightman, 2019). The $\xi_{V}$ and $\xi_{Z}$ are Gaussian noise with amplitude of 0.001 and 0.1. In most cases, the autoencoder and latent diffusion training can be completed within 1 to 3 hours on a single Nvidia A100 40G GPU.
Inference details.
We synthesize 100 novel parameters by feeding random noise into the latent diffusion model and the trained decoder. These synthesized parameters are then concatenated with the aforementioned fixed parameters to form our generated models. From these generated models, we select the one with the best performance on the training set. Subsequently, we evaluate its accuracy on the validation set and report the results. That is a consideration of making fair comparisons with the models trained using SGD optimization. We empirically find the performance on the training set is good for selecting models for testing.
Baselines.
1) The best validation accuracy among the original models is denoted as ‘original’. 2) Average weight ensemble (Krogh & Vedelsby, 1994; Wortsman et al., 2022) of original models is denoted as ‘ensemble’.
Table 2: p-diff main ablation experiments. We ablate the number of original models $K$, the location of applying our approach, and the effect of noise augmentation. The default settings are $K=200$, applying p-diff on the deep BN parameters (between layer16 to 18), and using noise augmentation in the input parameters and latent representations. Defaults are marked in gray. Bold entries are best results.
@@ 3.1 Setup
(a) Large $K$ can improve the performance stability of our method.
@@
(b) P-diff works well on deep layers. The index of layer is aligned with the standard ResNet-18.
@@
(c) Noise augmentation makes p-diff stronger. Adding noise on latent representations is more important than on parameters.
[…]
original ensemble p-diff parameter number
CIFAR-10 77.2 77.3 77.5 24714
CIFAR-100 57.2 57.2 57.3 70884
(d) Result comparisons on ConvNet-3 (includes three convolutional layers and one linear layer.
[…]
original ensemble p-diff parameter number
MNIST 85.3 85.2 85.4 39760
CIFAR-10 48.1 48.1 48.2 155135
(e) Result comparisons on MLP-3 (includes three linear layers and ReLU activation function).
3.2 Results
@@ 3.3 Ablation studies and analysis
Extensive ablation studies are conducted in this section to illustrate the characteristics of our method. We default to training ResNet-18 on CIFAR-100 and report the best, average, and medium accuracy (if not otherwise stated).
@@
Tab. 3(a) varies the size of training data, i.e. the number of original models. We find the performance gap of best results among different numbers of the original models is minor. To comprehensively explore the influences of different numbers of training data on the performance stability, we also report the average (avg.) and median (med.) accuracy as metrics of stability of our generated models. Notably, the stability of models generated with a small number of training instances is much worse than that observed in larger settings. This can be explained by the learning principle of the diffusion model: the diffusion process may be hard to model the target distribution well if only a few input samples are used for training.
@@
We default to synthesizing the parameters of the last two normalization layers. To investigate the effectiveness of p-diff on other depths of normalization layers, we also explore the performance of synthesizing the other shallow-layer parameters. To keep an equal number of BN parameters, we implement our approach to three sets of BN layers, which are between layers with different depths. As shown in Tab. 3(b), we empirically find that our approach achieves better performances (best accuracy) than the original models on all depths of BN layers settings. Another finding is that synthesizing the deep layers can achieve better accuracy than generating the shallow ones. This is because generating shallow-layer parameters is more likely to accumulate errors during the forward propagation than generating deep-layer parameters.
@@
Noise augmentation is designed to enhance the robustness and generalization of training the autoencoder. We ablate the effectiveness of applying this augmentation in the input parameters and latent representations, respectively. The ablation results are presented in Tab. 3(c). Several observations can be summarized as follows: i) Noise augmentation plays a crucial role in generating stable and high-performing models. ii) The performance gains of applying noise augmentation in the latent representations are larger than in the input parameters. iii) Our default setting, jointly using noise augmentation in parameters and representations obtains the best performances (includes best, average, and medium accuracy).
Generalization on entire model parameters.
[…]
We present result comparisons between our approach and two baselines (i.e., original and ensemble) at Tab. 3(e). We report the comparisons and parameter numbers of ConvNet-3 on CIFAR-10/100 and MLP-3 on CIFAR-10 and MNIST datasets. These experiments demonstrate the effectiveness and generalization of our approach in synthesizing entire model parameters, i.e., achieving similar or even improved performances over baselines. These results suggest the practical applicability of our method. However, we can not synthesize the entire parameters of large architectures, such as ResNet, ViT, and ConvNeXt series. It is mainly constrained by the limitation of the GPU memory.
Parameter patterns of original models.

### v2

# Neural Network Parameter Diffusion

@@ Abstract
Diffusion models have achieved remarkable success in image and video generation. In this work, we demonstrate that diffusion models can also generate high-performing neural network parameters. Our approach is simple, utilizing an autoencoder and a standard latent diffusion model. The autoencoder extracts latent representations of a subset of the trained network parameters. A diffusion model is then trained to synthesize these latent parameter representations from random noise. It then generates new representations that are passed through the autoencoder’s decoder, whose outputs are ready to use as new subsets of network parameters. Across various architectures and datasets, our diffusion process consistently generates models of comparable or improved performance over trained networks, with minimal additional cost. Notably, we empirically find that the generated models are not memorizing the trained networks. Our results encourage more exploration on the versatile use of diffusion models.
@@ 1 Introduction
The origin of diffusion models can be traced back to non-equilibrium thermodynamics [30, 65]. Diffusion processes were first utilized to progressively remove noise from inputs and generate clear images in [65]. Later works, such as DDPM [28] and DDIM [67], refine diffusion models, with a training paradigm characterized by forward and reverse processes.
@@
Despite the great success of diffusion models in visual generation, their potential in other domains remains relatively underexplored. In this work, we demonstrate the surprising capability of diffusion models in generating high-performing model parameters, a task fundamentally distinct from traditional visual generation. Parameter generation focuses on creating neural network parameters that can perform well on given tasks. It has been explored from prior and probability modeling aspects, i.e. stochastic neural network [66, 5, 74, 62, 46] and Bayesian neural network [48, 32, 58, 33, 20]. However, using a diffusion model in parameter generation has not been well-explored yet.
Figure 1: The left: illustrates the standard diffusion process in image generation. The right: denotes the parameter distribution of batch normalization (BN) during the training CIFAR-100 with ResNet-18. The upper half of the bracket: BN weights. The lower half of the bracket: BN biases.
Taking a closer look at the neural network training and diffusion models, diffusion-based image generation shares commonalities with the stochastic gradient descent (SGD) learning process in the following aspects (illustrated in Fig. 1): i) Both neural network training and the reverse process of diffusion models can be regarded as transitions from random noise/initialization to specific distributions. ii) High-quality images and high-performing parameters can also be degraded into simple distributions, such as Gaussian distribution, through multiple noise additions.
Based on the observations above, we introduce a novel approach for parameter generation, named Neural Network Parameter-Diffusion (p-diff), which employs a standard latent diffusion model to synthesize a new set of parameters. That is motivated by the fact that the diffusion model has the capability to transform a given random distribution to a specific one. Our method is simple, comprising an autoencoder and a standard latent diffusion model to learn the distribution of high-performing parameters. First, for a subset of parameters of models trained by the SGD optimizer, the autoencoder is trained to extract the latent representations for these parameters. Then, we leverage a standard latent diffusion model to synthesize latent representations from random noise. Finally, the synthesized latent representations are passed through the trained autoencoder’s decoder to yield new high-performing model parameters.
@@
2 Neural Network Parameter Diffusion
@@ 2.1 Preliminaries of diffusion models
Diffusion models typically consist of forward and reverse processes in a multi-step chain indexed by timesteps. We introduce these two processes in the following section:
@@
In contrast to the forward process, the reverse process aims to train a denoising network to recursively remove the noise from $x_{t}$. It moves backward on the multi-step chain as t decreases from $T$ to 0. Mathematically, the reverse process can be formulated as follows,
@@
$L_{\mathrm{dm}}=\mathcal{D}_{\text{KL}}(q(x_{t-1}|x_{t},x_{0})||p_{\theta}(x_{t-1}|x_{t})),$ (3)
where the $\mathcal{D}_{\text{KL}}(\cdot||\cdot)$ denotes the Kullback–Leibler (KL) divergence that is normally used to compute the difference between two distributions.
@@
The goal of training the diffusion model is to find the reverse transitions that maximize the likelihood of the forward transitions in each time step $t$. In practice, training equivalently consists of minimizing the variational upper bound. The inference procedure aims to generate novel samples from random noise via the optimized denoising parameters $\theta^{*}$ and the multi-step chains in the reverse process.
Figure 2: Our approach consists of two processes: parameter autoencoder and generation. Parameter autoencoder aims to extract the latent representations and reconstruct model parameters via the decoder. The extracted representations are used to train a standard latent diffusion model (LDM). In inference, the random noise is fed into LDM and trained decoder to generate parameters.
@@ 2.2 Overview
We propose neural network parameter diffusion (p-diff), which aims to generate high-performing parameters from random noise. As illustrated in Fig. 2, our method consists of two processes, named parameter autoencoder and generation. Given a set of trained high-performing models, we first select a subset of these parameters and flatten them into 1-dimensional vectors. Subsequently, we introduce an encoder to extract latent representations from these vectors, accompanied by a decoder responsible for reconstructing the parameters from latent representations. Then, a standard latent diffusion model is trained to synthesize latent representations from random noise. After training, we utilize p-diff to generate new parameters via the following chain: random noise $\rightarrow$ reverse process $\rightarrow$ trained decoder $\rightarrow$ generated parameters.
@@ 2.3 Parameter autoencoder
In our paper, we default to synthesizing a subset of model parameters. Therefore, to collect the training data for the autoencoder, we train a model from scratch and densely save checkpoints in the last epoch. It is worth noting that we only update the selected subset of parameters via SGD optimizer and fix the remained parameters of the model. The saved subsets of parameters $S=[s_{1},\ldots,s_{k},\ldots,s_{K}]$ are utilized to train the autoencoder, where $K$ is the number of the training samples. For some large architectures that have been trained on large-scale datasets, considering the cost of training them from scratch, we fine-tune a subset of the parameters of the pre-trained model and densely save the fine-tuned parameters as training samples.
@@
$L_{\mathrm{MSE}}=\frac{1}{K}\sum\begin{subarray}{c}K\\ \\ k=1\end{subarray}\|v_{k}-v_{k}^{\prime}\|^{2},$ (5)
where $v_{k}^{\prime}$ is the reconstructed parameters of the $k$-th model.
2.4 Parameter generation
One of the most direct strategies is to synthesize the novel parameters via a diffusion model. However, the memory cost of this operation is too heavy, especially when the dimension of $V$ is ultra-large. Based on this consideration, we apply the diffusion process to the latent representations by default. For $Z=[z_{1}^{0},\cdots,z_{k}^{0},\cdots,z_{K}^{0}]$ extracted from parameter autoencoder, we use the optimization of DDPM [28] as follows,
$\theta\leftarrow\theta-\nabla_{\theta}||\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}z_{k}^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)||^{2},$ (6)
where $t$ is uniform between $1$ and $T$, the sequence of hyperparameters $\bar{\alpha}_{t}$ indicates the noise strength at each step, $\epsilon$ is the added Gaussian noise, $\epsilon_{\theta}(\cdot)$ denotes the denoising network that parameterized by $\theta$. After finishing the training of the parameter generation, we directly feed random noise into the reverse process and the trained decoder to generate a new set of high-performing parameters. These generated parameters are concatenated with the remained model parameters to form new models for evaluation. Neural network parameters and image pixels exhibit significant disparities in several key aspects, including data type, dimensions, range, and physical interpretation. In contrast to images, neural network parameters mostly have no spatial relevance, so we replace 2D convolutions with 1D convolutions in our parameter autoencoder and parameter generation processes.
3 Experiments
In this section, we first introduce our experimental setup before reporting our results and ablation studies.
3.1 Setup
Datasets and architectures.
We evaluate our approach across a wide range of datasets, including MNIST [39], CIFAR-10/100 [36], ImageNet-1K [12], STL-10 [10], Flowers [50], Pets [51], and F-101 [4] to study the effectiveness of our method. We mainly conduct experiments on ResNet-18/50 [25], ViT-Tiny/Base [17], and ConvNeXt-T/B [42].
Training details.
The autoencoder and latent diffusion model both include a 4-layer 1D CNNs-based encoder and decoder. We default to collecting 200 training data for all architectures. For ResNet-18/50, we train the models from scratch. In the last epoch, we continue to train the last two normalization layers and fix the other parameters. We save 200 checkpoints in the last epoch, i.e., original models. For ViT-Tiny/Base and ConvNeXt-T/B, we fine-tune the last two normalization parameters of the released model in the timm library [73]. The $\xi_{V}$ and $\xi_{Z}$ are Gaussian noise with amplitude of 0.001 and 0.1. In most cases, the autoencoder and latent diffusion training can be completed within 1 to 3 hours on a single Nvidia A100 40G GPU.
Inference details.
We synthesize 100 sets of novel parameters by feeding random noise into the latent diffusion model and the trained decoder. These synthesized parameters are then concatenated with the aforementioned fixed parameters to form our generated models. From these generated models, we select the one with the best performance on the training set. Subsequently, we evaluate its accuracy on the validation set and report the results. That is a consideration of making fair comparisons with the models trained using SGD optimization. We empirically find the performance on the training set is good for selecting models for testing.
Baselines.
1) The best validation accuracy among the original models is denoted as ‘original’. 2) Average weight ensemble [38, 75] of original models is denoted as ‘ensemble’.
Table 1: We present results in the format of ‘original / ensemble / p-diff’. Our method obtains similar or higher performance than baselines. Our results are average in 3 runs. Bold entries are best results.
@@ 3.1 Setup
3.2 Results
Performance on partial parameter generation.
@@ 3.2 Results
original ensemble p-diff parameter number
CIFAR-10 77.2 77.3 77.5 24714
CIFAR-100 57.2 57.2 57.3 70884
(a) Result comparisons on ConvNet-3 (includes three convolutional layers and one linear layer).
[…]
original ensemble p-diff parameter number
MNIST 85.3 85.2 85.4 39760
CIFAR-10 48.1 48.1 48.2 155135
(b) Result comparisons on MLP-3 (includes three linear layers and ReLU activation function).
Generalization on entire parameter generation.
[…]
We present result comparisons between our approach and two baselines (i.e., original and ensemble) at Tab. 3(b). We report the comparisons and parameter numbers of ConvNet-3 on CIFAR-10/100 and MLP-3 on CIFAR-10 and MNIST datasets. These experiments demonstrate the effectiveness and generalization of our approach in synthesizing entire model parameters, i.e., achieving similar or even improved performances over baselines. These results suggest the practical applicability of our method. However, we can not synthesize the entire parameters of large architectures, such as ResNet, ViT, and ConvNeXt series. It is mainly constrained by the limitation of the GPU memory.
Table 3: p-diff main ablation experiments. We ablate the number of original models $K$, the location of applying p-diff, and the effect of noise augmentation. The default settings are $K=200$, applying p-diff on the deep BN parameters (between layer16 to 18), and using noise augmentation in the input parameters and latent representations. Defaults are marked in gray. Bold entries are best results.
@@
(c) Large $K$ can improve the performance stability of our method.
@@
(d) P-diff works well on deep layers. The index of layer is aligned with the standard ResNet-18.
@@
(e) Noise aug. makes p-diff stronger. Adding noise on latent representations is more important than on parameters.
@@ 3.3 Ablation studies and analysis
Extensive ablations are conducted to illustrate the characteristics of our method. We train ResNet-18 on CIFAR-100 and report the best, average, and medium accuracy (if not otherwise stated).
@@
Tab. 3(c) varies the size of training data, i.e. the number of original models. We find the performance gap of best results among different numbers of the original models is minor. To comprehensively explore the influences of different numbers of training data on the performance stability, we also report the average (avg.) and median (med.) accuracy as metrics of stability of our generated models. Notably, the stability of models generated with a small number of training instances is much worse than that observed in larger settings. This can be explained by the learning principle of the diffusion model: the diffusion process may be hard to model the target distribution well if only a few input samples are used for training.
@@
We default to synthesizing the parameters of the last two normalization layers. To investigate the effectiveness of p-diff on other depths of normalization layers, we also explore the performance of synthesizing the other shallow-layer parameters. To keep an equal number of BN parameters, we implement our approach to three sets of BN layers, which are between layers with different depths. As shown in Tab. 3(d), we empirically find that our approach achieves better performances (best accuracy) than the original models on all depths of BN layers settings. Another finding is that synthesizing the deep layers can achieve better accuracy than generating the shallow ones. This is because generating shallow-layer parameters is more likely to accumulate errors during the forward propagation than generating deep-layer parameters.
@@
Noise augmentation is designed to enhance the robustness and generalization of training the autoencoder. We ablate the effectiveness of applying this augmentation in the input parameters and latent representations, respectively. The ablation results are presented in Tab. 3(e). Several observations can be summarized as follows: i) Noise augmentation plays a crucial role in generating stable and high-performing models. ii) The performance gains of applying noise augmentation in the latent representations are larger than in the input parameters. iii) Our default setting, jointly using noise augmentation in parameters and representations obtains the best performances (includes best, average, and medium accuracy).
Table 4: Results on Adam and AdamW.
optimizer original best avg. med.
Adam [31] 70.5 70.5 64.6 69.3
AdamW [44] 71.5 71.6 67.0 70.5
Generalization to other optimizers.
We default to evaluating the effectiveness of our approach with SGD-trained models. To verify the robustness of p-diff with other optimizers, we evaluate the generalization of our method with Adam [31] and AdamW [44]. We follow the same data preparation, training, and evaluation process as claimed in the previous section. Here, we implement our method with ResNet-18 on the CIFAR-100 dataset. As shown in Tab. 4, our approach achieves comparable results to the original models, demonstrating good generalization with both Adam and AdamW.
(f) Similarity comparisons of original and p-diff models.
(g) Similarity comparisons of fine-tuned, noise-added, and p-diff models.
(h) t-SNE visualizations of original, p-diff, and adding noise.
Figure 3: The similarity represents the Intersection of Union (IoU) over wrong predictions between/among two models. (a) shows the comparisons in four cases: similarity among original models and p-diff models, similarity between original and p-diff models, and the maximum similarity (nearest neighbor) between original and p-diff models. (b) displays the accuracy and max similarity of fine-tuned, noise-added, and p-diff models. All the maximum similarities are calculated with the original models. (c) presents the t-SNE [71] of latent representations of the original models, p-diff models, and adding noise operation.